3.9Consider the following IS-LM model: C = Co + c1(Y-T)
I=b0 +b1Y-b2i
M/P = d1Y – d2i
a. Solve for equilibrium output. Assume (Assume c1 + b1 < 1). Now let investment depend on both sales and the interest rate: b. Solve for equilibrium level of interest rate.
Let’s go behind the scene in the monetary market. Use the equilibrium in the money market M/P = d1Y – d2i to solve for the equilibrium level of the real money supply.
How does the real money supply vary with government spending?
1
Expert's answer
2020-04-30T10:05:18-0400

The IS curve shifts to the left. Output falls at the same interest rate. Investment, which

depends positively on the level of output and negatively on the interest rate, also falls.

"Y=1\/(1-b1-c1)*(c0-c1T+b0-b2i+G)"

"I=b0+b1Y-b2i"

"I=b0+b1\/(1-b1-c3)*(c0-c1T+b0-b2i+G)+b2i"

"M\/P=d1Y-d2i"

"d*(M\/P)\/dG=d1\/(1-b1-c1)"